OpenStudy (lena772):

Determine the osmotic pressure (in atm) , at 80.2 o F, of an aqueous Iron (II) nitrate solution whose mole fraction of solute is 0.002696. Density solution = 1.11 g/mL

OpenStudy (lena772):

PV= nRT So Posm = nRT/V R= 0.08206 L*atm/mol*k T= 80.2F = 299.928 K V = calculate somehow using density n = reverse mole fraction calculation to find total moles of solution

OpenStudy (lena772):

@taramgrant0543664 @photon336 Can you guys assist me in reversing the mole fraction equation to find total moles of solution and then using that calculation with the density to find volume?

OpenStudy (taramgrant0543664):

V=m/d where v is volume, m is mass and d is density

OpenStudy (lena772):

right, but i can't find the mass of the solution until i reverse the mole fraction and find moles of total solution

OpenStudy (lena772):

I think? Mole fraction = (moles solute) / (moles solute + moles solv ). We are given moles of solute, but I don't know where to go from here. :(

OpenStudy (taramgrant0543664):

Yes that is the formula

OpenStudy (taramgrant0543664):

You're not given moles of solvent?

OpenStudy (lena772):

@Abhisar @pooja195 @zepdrix @abb0t

OpenStudy (lena772):

Nope. :/

OpenStudy (taramgrant0543664):

Hmmm... I don't know to tell you sorry I've only touched on osmotic pressure stuff none of my profs or teachers really talked much about it so I don't want to lead you on the wrong direction

OpenStudy (lena772):

Do you know anyone online that can help me? @taramgrant0543664

OpenStudy (taramgrant0543664):

I think you got a majority of the ones that could hopefully help

OpenStudy (lena772):

@abb0t Do you think you can help me?

OpenStudy (lena772):

@NoelGreco Hi! Can you please help me?

OpenStudy (mayankdevnani):

@Lena772 can you find out molality ?

OpenStudy (lena772):

yes moles solute/kg solvent. But I don't know the moles of solute.

OpenStudy (mayankdevnani):

Here is a formula :- \[\large \bf Molality=\frac{X_{solute}\times 1000 }{(1-X_{solute}) \times M_{solvent}}\]

OpenStudy (mayankdevnani):

and then use this :- \[\large \bf Molarity=\frac{molality \times density \times 1000}{molality \times Molar~ mass~ of~ solute+1000}\]

OpenStudy (mayankdevnani):

hope you can find molarity

OpenStudy (lena772):

Molal = (0.002696 mol Solute * 1000) / ( 0.997304 * molality of solvent) , BUT idk the molality of the solvent

OpenStudy (mayankdevnani):

its molar mass of solvent

OpenStudy (mayankdevnani):

not molality of the solvent Remember molality is denoted by small m i.e `m`

OpenStudy (lena772):

Molal = (0.002696 mol Solute * 1000) / ( 0.997304 * 175.8948)

OpenStudy (mayankdevnani):

after you find out molarity ! Use this :- \[\large \bf Osmotic~Pressure(\pi)=Molarity \times RT \times vant~hoff~factor\]

OpenStudy (lena772):

0.015368778 molal ?

OpenStudy (mayankdevnani):

i don't know. Use calculator

OpenStudy (mayankdevnani):

i am here to give you HINTS as possible

OpenStudy (lena772):

molality = 0.94396197 0.94396197 * 0.08206 * 299.928 K * van hoult?

OpenStudy (mayankdevnani):

you don't know vant hoff factor ?

OpenStudy (mayankdevnani):

denoted by `i`

OpenStudy (mayankdevnani):

@Lena772 then first learn this term and then solve the rest of question

OpenStudy (mayankdevnani):

any problem ???

OpenStudy (lena772):

Would the factor be 4 because there is 3 pieces of iron nitrate and 1 piece of water molecule?

OpenStudy (lena772):

@mayankdevnani

OpenStudy (mayankdevnani):

nope ! try again

OpenStudy (lena772):

What's wrong about it?

OpenStudy (mayankdevnani):

see the oxidation state of iron in question. You took the oxidation state as +3

OpenStudy (lena772):

So van hoff is 3 ?

OpenStudy (mayankdevnani):

yup

OpenStudy (lena772):

I get 69.7 atm. That's sounds extreme. Can you just plug those in and see if you get the same value?

OpenStudy (mayankdevnani):

lol but i don't have much time to solve ! Sorry ! Better luck ,Next time

OpenStudy (mayankdevnani):

i have to go now ! Bye ! @Lena772 Its been a nice time with you ! :)

OpenStudy (mayankdevnani):

MEDAL for your HARDWORK and keen interest on solving this question !

OpenStudy (mayankdevnani):

REMEMBER :- Don't suppose that i am a professor, i am (18)-